OpenStudy (anonymous):

An equation is shown below: 4x + 2(x – 3) = 4x + 2x – 11 Part A: Solve the equation and write the number of solutions. Show all the steps. Part B: Name one property you used to solve this equation.

OpenStudy (kidrah69):

I thought u said u got this?

OpenStudy (kidrah69):

Anyways 2(x – 3) distribute

OpenStudy (kidrah69):

Just how i said in the previous problem.

OpenStudy (anonymous):

okay... one second.

OpenStudy (anonymous):

4x + 2x - 6 = 4x + 2x - 11

OpenStudy (anonymous):

???

OpenStudy (kidrah69):

ok now combine all the like terms to isolate X

OpenStudy (anonymous):

wait what? like this? 6x - 6 = 6x - 11

OpenStudy (kidrah69):

good now retrice6 to both sides

OpenStudy (kidrah69):

ONLY 6 not 6x

OpenStudy (anonymous):

you would have to divide 6 by -11? or -11 by 6?

OpenStudy (anonymous):

wait no...

OpenStudy (kidrah69):

nope just ADD 6

OpenStudy (anonymous):

-6 = -11

OpenStudy (kidrah69):

-11+6=?

OpenStudy (anonymous):

theres no solutions..

OpenStudy (anonymous):

but what do they mean for me to name a property?

OpenStudy (kidrah69):

Ok part a is right by property what did u use to solve it we distributed so what property?

OpenStudy (anonymous):

idk... :(

OpenStudy (kidrah69):

Which one ? :)

OpenStudy (anonymous):

the first one...?

OpenStudy (kidrah69):

we multiplied...

OpenStudy (kidrah69):

it cant be the 1st

OpenStudy (anonymous):

the 4th or 5th?

OpenStudy (anonymous):

im so sorry that i look so stupid. lol

OpenStudy (kidrah69):

dun call yrself stupid :P yr smart u got it! its the 5th one :)

OpenStudy (anonymous):

oh okay, well thank you so so much for all your help. i really appreciate it! :)

OpenStudy (kidrah69):

no probs :)